2016 2017 $ 65,058 32,535 18,568 $ 65,299 32,909 18,949 13,441 579 182 Amounts in millions except per share amounts; Years en2017 $ 15,411 2016 $ 10,604 2015 $ 7,144 (1,679) 1 Amounts in millions; Years ended June 30 NET EARNINGS OTHER COMPREHENSIVE$ $ 9,632 7,024 13,554 30,210 18,038 8,126 8,254 64,628 9,325 7,449 2,343 11,653 30,770 18,945 9,113 10,325 69,153 LiabilitiePaid-In Capital $63,911 ESOP Debt Retirement ($1,340) Income Treasury Retained controlling holders (Loss) Stock Earnings Int(6,989) (247) (6,989) (247) (14,625) 2,981 Common Preferred, net of tax benefits Treasury purchases (2) Employee plan issuanc(149) (43) 162 1,285 204 184 15,435 928 (976) 746 14,608 12,753 (3,384) 571 (3,314) 432 (3,736) 4,498 (908) (137) (3,647) 1,2TUUUUUU IIVIN NUN WILL TITULLUTTUUN VIIVIL LUI IYUNLIITUIT | III (4,931)(1) 1940 1,34 Pre-divestiture addition of restrictedP&Gs net cash provided by operating activities ✓ by % from 2016 to 2017. e Textbook and Media Compute P&Gs (1) current cashprovided by operating activities over the period 2015-2017? (Enter negative amounts using either a negative sign preceding thWhat were P&Gs cash flows from its operating, investing, and financing activities for 2017? What were its trends in net cashHow much working capital did P&G have on June 30, 2017? On June 30, 2016? (Enter negative amounts using either a negative sig
Answer #1

1)

  • CURRENT CASH DEBT COVERAGE RATIO (2017)
    = Operating Cash Flows / Average of current liabilities
    = 12,753 / (30,210/30,770)2
    = 12,753 / 30,490
    = 0.418 times.
  • CASH DEBT COVERAGE RATIO (2017)
    = Opertaing Cash Flows / Total Liabilities
    = 12,753 / 30,210
    = 0.422 times.
  • FREE CASH FLOW (2017)
    = Cash flow from Operating Activities - Capital Expenditure Incurred
    = 12,753 - 3,384
    = 9,369

2)

Particulars Working 2017 2016
cash provided in operating activities Given 12,753 15,435
cash used in investing activities Given 5,689 5,575
cash used in financing activities Given 8,568 9,213

3)

Particulars Amount %
Net Cash provided by operating Activities (2015-16) 14,608 - 15,435 = 827 5.67%
(827/14608)*100
Net Cash provided by operating Activities (2016-17) 15,435 - 12,753 = (2682) -17.37%
((2682) / 15435)*100

4)

Particulars June 30, 2017 June 30, 2016
Working Capital (Current Assets - Current Liabilities) 26,494 - 30,210 = -3,716 33,782 - 30,770 = 3,012